# Multi-Objective min-max Online Convex Optimization

### v1

@@ Appendix A Proof of Lemma 5
Appendix B Proof of Lemma 6
@@
Appendix C Proof of Lemma 7
@@
$R_{3}$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\min_{x\in\mathcal{X}}\max_{\theta}\sum_{t=1}^{T}\theta^{\top}\lambda_{t}(x)$
$\stackrel{{\scriptstyle(a)}}{{=}}\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-\sum_{t=1}^{T}\theta^{\star}{}^{\top}\lambda_{t}(x^{\star}),$
$=\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star}),$ (25)
where $(a)$ follows from the definition of $x^{\star}$ and $(b)$ follows since $x^{\star}\in\mathcal{X}$ is feasible. Moreover, we have
$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}{\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}),$ (26)
where the first equality follows since $\theta_{t}$ is completely determined by $\mathcal{F}_{t-1}$, while the second equality follows because ${\mathbb{E}}\{\lambda_{t}(x^{\star})|\mathcal{F}_{t-1}\}=\mu(x^{\star})$ from the i.i.d. assumption.
@@
for each $t=1,\dots,T$ using the optimality of $\theta^{\star}$ (24) since
$\theta^{\star}\in\arg\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x^{\star})$
depends only on the expected vector $\mu(x^{\star})$ and not on any particular realization of $\lambda_{t}(x^{\star})$.
Thus, we get from (26)
$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}\leq 0,$ (27)
Using (27) in (25) and taking the expectation (using the tower property of conditional expectation),
$\mathbb{E}\{R_{3}\}\leq\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star})\leq 0.$ (28)
Appendix D Per-slot $\it{\it min-max}$} multi-objective optimization [24]
@@
Appendix E Experts Problem (Gains Version)

### v3

@@ Appendix E Proof of Lemma 6
Appendix F Proof of Lemma 7
@@
Appendix G Proof of Lemma 8
@@
$R_{3}$ $=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-T\min_{x\in\mathcal{X}}\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x)$
$=\min_{x\in\mathcal{X}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x)-T\,\theta^{\star}{}^{\top}\mu(x^{\star}),$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star}),$ (21)
where $(b)$ follows since $x^{\star}\in\mathcal{X}$ is feasible.
Now rewrite the right-hand side of (21) as
$\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star})$ $=\sum_{t=1}^{T}(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\;+\;\sum_{t=1}^{T}\theta^{\star}{}^{\top}\big(\lambda_{t}(x^{\star})-\mu(x^{\star})\big).$ (22)
Moreover, we have
$\mathbb{E}\left\{(\theta_{t}-\theta^{\star})^{\top}\lambda_{t}(x^{\star})\,\Big|\,\mathcal{F}_{t-1}\right\}=(\theta_{t}-\theta^{\star})^{\top}{\mathbb{E}}\{\lambda_{t}(x^{\star})\mid\mathcal{F}_{t-1}\}=(\theta_{t}-\theta^{\star})^{\top}\mu(x^{\star}),$ (23)
where the first equality follows since $\theta_{t}$ is completely determined by $\mathcal{F}_{t-1}$, while the second equality follows from the i.i.d. assumption.
@@
for each $t=1,\dots,T$, by the optimality of $\theta^{\star}$ in
$\theta^{\star}\in\arg\max_{\theta\in\Delta_{2}}\theta^{\top}\mu(x^{\star}).$
Further, note that
${\mathbb{E}}\!\left\{\theta^{\star}{}^{\top}\big(\lambda_{t}(x^{\star})-\mu(x^{\star})\big)\,\Big|\,\mathcal{F}_{t-1}\right\}=0,$
again by the i.i.d. assumption.
Combining the above and taking conditional expectations in (22), we obtain
${\mathbb{E}}\!\left\{\sum_{t=1}^{T}\theta_{t}^{\top}\lambda_{t}(x^{\star})-T\,\theta^{\star}{}^{\top}\mu(x^{\star})\right\}\leq 0.$ (24)
Hence,
$\mathbb{E}\{R_{3}\}\leq 0.$ (25)
Appendix H Per-slot $\it{\it min-max}$} multi-objective optimization [25]
@@
Remark 11.
Detailed comparison with $\alpha$-fair model of [40] The transition from $\alpha$-fairness (with finite $\alpha$) (6) to the min-max objective (1) marks a shift from a smooth, strictly concave optimization landscape to a non-smooth, polyhedral one. For any finite $\alpha\geq 0$, the $\alpha$-fair utility function $U_{\alpha}(z)$ (6) acts as a differentiable, strictly concave regularizer that “softens” the trade-offs between competing objectives. In the finite $\alpha$ regime, the gradient of the aggregate utility changes continuously with the input, allowing algorithms to rely on the local curvature of the utility function to guide the learning process. However, as $\alpha\to\infty$, the utility function converges to the min-max objective, which is equivalent to an $L_{\infty}$ norm. Geometrically, the smooth, rounded contours of the $\alpha$-fair utility collapse into a sharp wedge where the derivative is discontinuous at the points where multiple objectives have equal costs. This loss of differentiability explains why prior regret bounds of $O(\alpha\sqrt{T})$ [40] become vacuous in the min-max limit; they are inversely proportional to the curvature of the utility, which vanishes as the objective becomes piecewise linear.
In comparison, the Hedge+OGD algorithmic approach succeeds as $\alpha\rightarrow\infty$ because it does not attempt to “smoothen” this non-differentiable cusp. Instead, by employing the Hedge algorithm to update weights $\theta_{t}$, it treats the selection of the dominant objective as a discrete experts problem. This effectively bypasses the numerical instability inherent in the $\alpha\to\infty$ transition and provides a regret guarantee that remains independent of the parameter $\alpha$.
Hedge+OGD algorithm is also applicable for finite $\alpha$, however, since it is designed to remain stable in the non-smooth limit ($\alpha\to\infty$), it lacks the specialized mechanics required to exploit the curvature present when $\alpha$ is small. For finite $\alpha$, the $\alpha$-fair utility function is strictly concave, offering a smooth landscape for optimization. The OHF algorithm [40] is specifically tailored to exploit this strict concavity, allowing the algorithm to achieve a regret bound of $O(\alpha\sqrt{T})$. When $\alpha$ is small, the utility function possesses significant curvature, which provides a strong, directional gradient signal that accelerates convergence. Hedge on the other hand is a first-order expert algorithm designed for the simplex; it does not naturally incorporate the second-order information or the strong concavity of the underlying utility. Consequently, it cannot “speed up” when the landscape is smooth.
The main takeaway is that the finite $\alpha$ and $\alpha\rightarrow\infty$ regimes are fundamentally different and require different treatment.
Appendix I Proof of Theorem 9
Recall the decomposition (10),
Bounding $R_{1}$ and $R_{3}$.
Recall that
$R_{1}:=\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t}),\qquad\Lambda_{t}(x,\theta)=\sum_{k=1}^{K}\theta_{k}f_{t}^{k}(x).$
In the bandit setting, the Hedge update observes the queried losses $\{f_{t-1}^{k}(\tilde{x}_{t})\}_{k=1}^{K}$. To relate $R_{1}$ to the Hedge regret on the queried points, we proceed as follows.
Step 1 (add and subtract queried losses inside the maximization). For any fixed $\theta\in\Delta_{K}$,
$\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta)=\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)+\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr).$
Substituting this identity into the definition of $R_{1}$ yields
$R_{1}$ $=\max_{\theta\in\Delta_{K}}\Bigg[\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$
$\hskip 85.35826pt+\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr)\Bigg].$
Step 2 (split the maximization). Using the elementary inequality
$\max_{\theta}(a_{\theta}+b_{\theta})\;\leq\;\max_{\theta}a_{\theta}+\max_{\theta}b_{\theta},$
we obtain
$R_{1}$ $\leq\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})$
$\quad+\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr).$
Step 3 (Lipschitz bound, uniform in $\theta$). By Assumption 3, each $f_{t}^{k}$ is $G$-Lipschitz, and since $\theta\in\Delta_{K}$ is a convex combination,
$\bigl|\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr|\leq G\|x_{t}-\tilde{x}_{t}\|,\qquad\forall\,\theta\in\Delta_{K}.$
Therefore,
$\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\bigl(\Lambda_{t}(x_{t},\theta)-\Lambda_{t}(\tilde{x}_{t},\theta)\bigr)\leq G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
Combining the above bounds gives
$R_{1}\leq\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})+G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
Step 4 (remove the remaining mismatch outside the maximization). Adding and subtracting $\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})$ yields
$R_{1}$ $\leq\Bigg[\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\Bigg]$
$\quad+\Bigg[\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})-\sum_{t=1}^{T}\Lambda_{t}(x_{t},\theta_{t})\Bigg]+G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
The second bracket is again bounded by Lipschitz continuity:
$\bigl|\Lambda_{t}(\tilde{x}_{t},\theta_{t})-\Lambda_{t}(x_{t},\theta_{t})\bigr|\leq G\|x_{t}-\tilde{x}_{t}\|.$
Final bound on $R_{1}$.
$R_{1}\leq\underbrace{\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})}_{\text{Hedge regret on queried losses}}\;+\;2G\sum_{t=1}^{T}\|x_{t}-\tilde{x}_{t}\|.$
The first term is exactly the regret of the $K$-experts problem with gains $f_{t-1}^{k}(\tilde{x}_{t})$, and hence with $\eta_{\theta,t}=\sqrt{\log K}/(B\sqrt{t})$,
$\max_{\theta\in\Delta_{K}}\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta)-\sum_{t=1}^{T}\Lambda_{t}(\tilde{x}_{t},\theta_{t})\leq B\sqrt{2T\log K}.$
Moreover, $\|x_{t}-\tilde{x}_{t}\|\leq\delta$, so the additional term $2G\sum_{t}\|x_{t}-\tilde{x}_{t}\|$ is of order $\sum_{t}\delta$. We upper bound this in the two cases as follows.
  • One-point feedback. With $\delta=T^{-1/4}$,
    $2G\sum_{t=1}^{T}\delta=O\!\left(G\sum_{t=1}^{T}T^{-1/4}\right)=O(GT^{3/4}).$
  • Two-point feedback. With $\delta=T^{-1/2}$,
    $2G\sum_{t=1}^{T}\delta=O\!\left(G\sum_{t=1}^{T}T^{-1/2}\right)=O(G\sqrt{T}).$
Note that the term $R_{3}$ depends only on the true loss functions $(f_{t},g_{t})$ and the sequence $\{\theta_{t}\}_{t=1}^{T}$, and not on how gradients are estimated; hence the proof of ${\mathbb{E}}\{R_{3}\}\leq 0$ from Appendix G applies verbatim in the bandit setting as well.
Thus,
${\mathbb{E}}\{R_{3}\}\leq 0.$
We bound $R_{2}$ as follows.
Bounding $R_{2}$ (one-point feedback). Using $\delta=T^{-1/4}$ and $\eta_{x,t}=D/(Gd^{1/2}t^{3/4})$, $\rho=\delta/{\mathsf{r}}$ where $r\mathcal{B}_{1}\subseteq\mathcal{X}\subseteq D\mathcal{B}_{1}$ and $\mathcal{B}_{1}$ is the unit ball, the bandit-OGD analysis [16] yields
$R_{2}\leq c_{1}\,d^{1/2}GD\,T^{3/4},$
for some universal constant $c_{1}$.
Combining the three bounds we get
${\bar{R}}_{\mathcal{A}}(T)\leq c_{1}\,d^{1/2}GD\,T^{3/4}\;+\;2B\sqrt{2T\log K}.$
Bounding $R_{2}$ (two-point feedback). Choosing $\delta=T^{-1/2}$, $\rho=\delta/{\mathsf{r}}$ and $\eta_{x,t}=D/(Gd^{1/2}\sqrt{t})$, the two-point bandit-OGD analysis [1, 13, 39] gives
$R_{2}\leq c_{3}\,d^{1/2}GD\,\sqrt{T},$
for some universal constant $c_{3}$. Thus, with two-point feedback,
${\bar{R}}_{\mathcal{A}}(T)\leq c_{3}\,d^{1/2}GD\,\sqrt{T}\;+\;2B\sqrt{2T\log K}.$
Appendix J Experts Problem (Gains Version)